### v1

@@ Abstract
An utterance-level speaker embedding is typically obtained by aggregating a sequence of frame-level representations. However, in real-world scenarios, individual frames encode not only speaker-relevant information but also various nuisance factors. As a result, different frames contribute unequally to the final utterance-level speaker representation for Automatic Speaker Verification systems. To address this issue, we propose to estimate the inherent uncertainty of each frame and assign adaptive weights accordingly, where frames with higher uncertainty receive lower attention. Based on this idea, we present $\mathcal{U}^{3}$-xi, a comprehensive framework designed to produce more reliable and interpretable uncertainty estimates for speaker embeddings. Specifically, we introduce several strategies for uncertainty supervision. First, we propose speaker-level uncertainty supervision via a Stochastic Variance Loss, where the distance between an utterance embedding and its corresponding speaker centroid serves as a pseudo ground truth for uncertainty learning. Second, we incorporate global-level uncertainty supervision by injecting the predicted uncertainty into the softmaxsoftmax scale during training. This adaptive scaling mechanism adjusts the sharpness of the decision boundary according to sample difficulty, providing global guidance. Third, we redesign the uncertainty estimation module by integrating a Transformer encoder with multi-view self-attention, enabling the model to capture rich local and long-range temporal dependencies. Comprehensive experiments demonstrate that $\mathcal{U}^{3}$-xi is model-agnostic and can be seamlessly applied to various speaker encoders. In particular, when applied to ECAPA-TDNN, it achieves 21.1% and 15.57% relative improvements on the VoxCeleb1 test sets in terms of EER and minDCF, respectively.
@@ II-B Large Margin Softmax Loss Function
Large margin softmax loss functions have been extensively studied and demonstrated to be highly effective in both face recognition [40, 41, 42, 43, 44, 45, 46] and speaker verification tasks [47, 48, 49, 50]. By introducing an angular or additive margin to the target class, these losses enforce stricter separation between classes and encourage intra-class compactness and inter-class discrimination. Early works such as SphereFace [41] introduced multiplicative angular margins, while subsequent methods, including CosFace (additive cosine margin) [42] and ArcFace (additive angular margin) [43], further improved the geometric interpretability and optimization stability of margin constraints. In the speaker verification domain, large margin softmax variants (e.g., AAM-Softmax [43], AM-Softmax [42]) have become standard training objectives due to their ability to produce highly discriminative speaker embeddings.
In addition to margin manipulation, several studies [51, 52, 53, 54, 55, 52, 56] have explored the role of the scale parameter, which controls the sharpness of the class posterior distribution produced by the output of the softmax layer. A larger scale magnifies the logit differences, yielding a sharper posterior distribution and stronger class separation. However, it may also cause over-confidence when the embeddings are unreliable. Conversely, a smaller scale smooths the posterior distribution and stabilizes gradients, especially under challenging acoustic conditions, though at the cost of reduced discriminability. This is because a smaller scale dampens the logit values sent to the softmax layer, making extreme probabilities (near 0 or 1) less common. With fewer extreme probabilities, the loss function’s gradients change more gradually during training, avoiding sharp fluctuations or vanishing gradients that could disrupt learning. Different works leverage this trade-off in distinct ways. It has been observed that using a smaller scale can enhance adversarial robustness [53], whereas AdaCos [54] adopts a dynamic scale that gradually decreases throughout training. In contrast, other works increase the scale dynamically, arguing that a larger scale facilitates learning of hard samples [55, 51]. Beyond fixed or manually designed schedules, some methods propose predicting a sample-wise scale using DNNs [52, 56]. In these approaches, the learned scale can be interpreted as a quality or uncertainty indicator, with higher predicted scales corresponding to cleaner or more easily classifiable samples, and lower scales reflecting higher uncertainty or degraded quality.
@@ II-C Temporal Pooling Layers
Temporal pooling plays a critical role in speaker embedding extraction. It aggregates sequences of frame-level features into utterance-level representations, from which the final speaker embeddings are derived [2, 57]. Existing temporal pooling approaches range from relatively simple statistical methods to more sophisticated techniques incorporating attention mechanisms. Statistical pooling methods summarize the frame sequence using global statistics. Temporal Average Pooling (TAP) [8] computes the mean of frame-level deep features to represent the utterance-level embedding. Temporal Statistics Pooling (TSTP) [9] extends TAP by concatenating the mean and standard deviation vectors, thereby incorporating variance information. Temporal Standard Deviation Pooling (TSDP) [10] demonstrates that even using only the standard deviation vector can remain effective.
However, traditional statistical pooling assumes that all frames contribute equally to the utterance-level representation, which is often not the case in practice. Frame-level features may vary in their speaker-discriminative importance, motivating the use of attention mechanisms to weight frames differently. Attentive Statistics Pooling (ASP) [11] addresses this issue by computing both weighted mean and weighted standard deviation vectors. Subsequent studies [12, 13, 14, 15] demonstrate that multi-head attention further improves pooling performance. ECAPA-TDNN [16] advances this idea by extending attention to the channel dimension, allowing the model to emphasize speaker characteristics that do not activate at the same temporal positions.
In the xi-vector model [24], attention-based weighting is replaced by a sequence of frame-level uncertainty estimates, where higher uncertainty corresponds to a lower contribution to the utterance-level representation. As adopted in this paper, this uncertainty-aware pooling strategy provides a principled and interpretable mechanism for down-weighting unreliable frames while preserving speaker-discriminative information.
@@ III-A Uncertainty-aware Speaker Model: xi-vector
$\boldsymbol{\phi}^{s}=\left(\frac{\boldsymbol{\phi}-\boldsymbol{\mu}_{\text{bn}}}{\sqrt{\boldsymbol{\sigma}_{\text{bn}}+\epsilon\mathbf{I}}}\otimes\boldsymbol{\gamma}_{\text{bn}}+\boldsymbol{\beta}_{\text{bn}}\right)\mathbf{A}_{\text{fc}}^{\top}+\mathbf{b}_{\text{fc}},$ (5)
@@
$\mathbf{\Sigma}^{s}=\mathbf{A}_{\text{fc}}\frac{\mathbf{L}^{-1}\otimes\boldsymbol{\gamma}_{\text{bn}}^{2}}{\boldsymbol{\sigma}_{\text{bn}}+\epsilon\mathbf{I}}\mathbf{A}_{\text{fc}}^{\top}.$ (6)
@@ III-B Softmax in Speaker Classification
As shown in Fig. 1, the projection layer maps the utterance-level representation $\boldsymbol{\phi}^{s}$ into a $C$-dimensional vector, where $C$ is the number of speakers in the training set. This vector is then optimized using a classification objective, with the softmax function playing a central role.
@@ III-B2 Additive Angular Margin Softmax
While the conventional Softmax loss serves as a basic classification objective, it does not explicitly encourage the learned embeddings to be both intra-class compact (features from the same class clustering tightly together) and inter-class separable (features from different classes being well apart). To overcome this limitation, the Additive Angular Margin Softmax (AAM-Softmax) loss [43] was proposed. This formulation strengthens discriminative power by introducing an angular margin penalty.
Specifically, AAM-Softmax first removes the bias term ($\mathbf{b}=0$) and reformulates the logit as $\mathbf{W}^{\top}\boldsymbol{\phi}^{s}=\|\mathbf{W}\|\|\boldsymbol{\phi}^{s}\|\cos\theta$, where $\theta$ is the angle between the class weight $\mathbf{W}$ and the feature embedding $\boldsymbol{\phi}^{s}$. Furthermore, to ensure that the prediction depends solely on the angular similarity rather than the vector magnitude, both $\mathbf{W}$ and $\boldsymbol{\phi}^{s}$ are normalized:
$\hat{\mathbf{W}}=\frac{\mathbf{W}}{\|\mathbf{W}\|},\ \hat{\boldsymbol{\phi}}^{s}=\frac{\boldsymbol{\phi}^{s}}{\|\boldsymbol{\phi}^{s}\|}.$ (8)
With these normalizations, the classification is determined by the cosine similarity. To further enhance discriminability, an additive angular margin $m$ is imposed on the target logit, leading to the following loss function:
$\mathcal{L}_{\text{AAM}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{s\cos(\theta_{y_{i}}+m)}}{e^{s\cos(\theta_{y_{i}}+m)}+\sum_{j=1,j\neq y_{i}}^{C}e^{s\cos\theta_{j}}},$ (9)
where $s$ is a fixed scaling factor that stabilizes optimization, and $m$ is the angular margin encouraging tighter intra-class clusters and larger inter-class separation.
@@ IV-A Stochastic Variance Loss
Our previous work xi-vector [24] relies solely on a classification loss to implicitly supervise frame-level uncertainty, without any form of explicit guidance. Such implicit learning often results in suboptimal or unstable uncertainty estimation. To overcome this limitation, we introduce an explicit supervision mechanism for utterance-level uncertainty modeling, termed Stochastic Variance Loss (SVL). Our previous work [36] firstly proposed SVL in the context of speaker recognition and demonstrated its effectiveness. In this paper, we further generalize this idea by exploring other alternative formulations to realize the core concept of SVL.
The proposed loss $\mathcal{L}_{\text{SVL}}$ aims to establish a pseudo ground truth for utterance-level uncertainty estimation. Specifically, we treat the distance between the centroid of embeddings from the same speaker and the current utterance’s embedding $\boldsymbol{\phi}^{s}$ as a pseudo label. Since only the embedding distance of a single utterance-centroid pair is used to approximate the uncertainty (i.e., variance), we refer to this approach as stochastic variance learning.
Suppose we use $\boldsymbol{\phi}^{c}_{y_{i}}$ to represent the centroid of speaker $y_{i}$ of utterance $i$:
$\boldsymbol{\phi}^{c}_{y_{i}}=\frac{\sum_{i=1}^{I}\boldsymbol{\phi}^{s}_{i}}{I},$ (10)
where all $I$ utterances belong to the same speaker $y_{i}$. The the $\mathcal{L}_{\text{SVL}}$ is given by:
$\mathcal{L}_{\text{SVL}}=\frac{1}{B}\sum_{b=1}^{B}\|\alpha\mathbf{(\Sigma}^{s}_{b})^{\frac{1}{2}}-|\boldsymbol{\phi}^{s}_{b}-\boldsymbol{\phi}^{c}_{y_{b}}|\|^{2},$ (11)
where $B$ denotes the number of batch size in training. $\mathbf{\Sigma}^{s}_{b}$ and $\boldsymbol{\phi}^{c}_{y_{b}}$ denote the covariance matrix for utterance $b$ and the centroid of speaker $y_{b}$, respectively. $\|\cdot\|$ denotes the Euclidean norm. In addition, $\alpha$ is a learnable parameter initialized from $1$, reference to as the uncertainty scaling factor, aiming to address potential scale mismatch between predicted uncertainty and pseudo ground truth.
@@
  • Pre: The centroid is computed using embeddings $\boldsymbol{\phi}^{s}$ extracted from a pre-trained model with the same architecture. But this pre-trained model is trained solely with the cross-entropy loss $\mathcal{L}_{\text{AAM}}$, without the proposed $\mathcal{L}_{\text{SVL}}$. As a result, the speaker centroids remain fixed throughout training and are subsequently used to provide supervision, as defined in Eq. (11).
  • Epoch: Instead of relying on a pre-trained model, the speaker embedding $\boldsymbol{\phi}^{s}$ is extracted using the checkpoint from the last epoch. Then then centroid is computed and updated after each training epoch. This allows the centroid representations to evolve along with model optimization, leading to progressively more accurate speaker representations.
@@
where $\kappa$ is used to control loss weight and is dynamically adjusted based on the current epoch $ep$ to ensure that $\mathcal{L}_{\text{SVL}}$ gradually contributes as training progresses:
$\kappa=\begin{cases}0,&\text{if }\text{epoch}\leq ep_{\text{SVL}},\\ \lambda\frac{ep-ep_{\text{SVL}}}{ep_{\text{Max}}-ep_{\text{SVL}}},&\text{otherwise},\end{cases}$ (13)
where $\lambda$ is a pre-defined fixed constant, $ep_{\text{SVL}}$ indicates epoch when $\mathcal{L}_{\text{SVL}}$ starts to be applied. $ep_{\text{Max}}$ indicates the maximum number of training epochs. In this paper, we set $ep_{\text{SVL}}=40$ and $ep_{\text{Max}}=150$. Cause we use WeSpeaker [58, 59] as training tool, its margin changes before $40$ epoch, which causes unstable training and produces NaN if we incorporate $\mathcal{L}_{\text{SVL}}$ at this stage.
@@ IV-B Uncertainty-aware Scoring
$s_{u}=\frac{<\boldsymbol{\phi}^{s}_{2},\boldsymbol{\phi}^{s}_{1}>}{\sqrt{\boldsymbol{\phi}{{}^{s}_{2}}^{\top}(\mathbf{I}+\rho\boldsymbol{\Sigma}^{s}_{2})^{-1}\boldsymbol{\phi}^{s}_{2}}\sqrt{\boldsymbol{{\phi}}{{}^{s}_{1}}^{\top}(\mathbf{I}+\rho\boldsymbol{\Sigma}^{s}_{1})^{-1}\boldsymbol{\phi}^{s}_{1}}},$ (14)
where $\rho$ denotes a scaling factor in cosine scoring, and $\mathbf{I}$ is the identity matrix. The subscripts $1$ and $2$ indicate the corresponding speaker identities for the enrollment and test embeddings, respectively. When the covariance matrix $\boldsymbol{\Sigma}^{s}$ or $\rho$ is zero, the uncertainty-aware cosine scoring $s_{u}$ reduces to the normal cosine similarity. Wang et al. [38] have proposed to use the fixed constant $1/d$ as $\rho$, where $d$ denotes the dimension of speaker embedding. However, a fixed value may not be optimal, as demonstrated in our previous work, xi+ [36]. To address this limitation, we introduce a learnable uncertainty scaling factor $\alpha$, which replaces $\rho$ in uncertainty-aware scoring. This parameter can be adapted to different training datasets, allowing the model to optimally calibrate the influence of uncertainty in the scoring process.

### v2

@@ Abstract
An utterance-level speaker embedding is typically obtained by aggregating a sequence of frame-level representations. However, in real-world scenarios, individual frames encode not only speaker-relevant information but also various nuisance factors. As a result, different frames contribute unequally to the final utterance-level speaker representation for Automatic Speaker Verification systems. To address this issue, we propose to estimate the inherent uncertainty of each frame and assign adaptive weights accordingly, where frames with higher uncertainty receive lower attention.Based on this idea, we present $\mathcal{U}^{3}$-xi (pronounced as “U-cube-xi”), a comprehensive framework designed to produce more reliable and interpretable uncertainty estimates for speaker embeddings. Specifically, we introduce several strategies for uncertainty supervision. First, we propose speaker-level uncertainty supervision via a Stochastic Variance Loss, where the distance between an utterance embedding and its corresponding speaker centroid serves as a pseudo ground truth for uncertainty learning. Second, we incorporate global-level uncertainty supervision by injecting the predicted uncertainty into the softmaxsoftmax scale during training. This adaptive scaling mechanism adjusts the sharpness of the decision boundary according to sample difficulty, providing global guidance. Third, we redesign the uncertainty estimation module by integrating a Transformer encoder with multi-view self-attention, enabling the model to capture rich local and long-range temporal dependencies. Comprehensive experiments demonstrate that $\mathcal{U}^{3}$-xi is model-agnostic and can be seamlessly applied to various speaker encoders. In particular, when applied to ECAPA-TDNN, it achieves 21.1% and 15.57% relative improvements on the VoxCeleb1 sets in terms of EER and minDCF, respectively  [^1] The code can be found in: https://github.com/mrjunjieli/wespeaker_u_cube.git.
@@ II-B Large Margin Softmax Loss Function
Large margin softmax loss functions have been extensively studied and shown to be highly effective in both face recognition [40, 41, 42, 43, 44, 45, 46] and speaker verification tasks [47, 48, 49, 50]. By introducing an angular or additive margin to the target class, these losses enforce stricter class separation and encourage intra-class compactness. Early works such as SphereFace [41] introduced multiplicative angular margins, while subsequent methods, including CosFace (additive cosine margin) [42] and ArcFace (additive angular margin) [43], further improved the geometric interpretability and optimization stability. In the speaker verification domain, large margin softmax variants (e.g., AAM-Softmax [43], AM-Softmax [42]) have become standard training objectives due to their strong discriminative capability.
In addition to margin design, several studies [51, 52, 53, 54, 55, 56] have investigated the role of the scale parameter, which controls the sharpness of the posterior distribution produced by the softmax layer. A larger scale amplifies logit differences, yielding sharper posteriors and stronger class separation, but may lead to over-confidence when embeddings are unreliable. Conversely, a smaller scale produces smoother posteriors and more stable gradients, particularly under challenging acoustic conditions, at the cost of reduced discriminability. Different works exploit this trade-off in various ways. For example, a smaller scale has been shown to improve adversarial robustness [53], while AdaCos [54] adopts a dynamic scale schedule. In contrast, other methods increase the scale during training to emphasize hard samples [55, 51]. Beyond manually designed strategies, several approaches predict sample-wise scales using DNNs [52, 56], where the learned scale can be interpreted as a quality or uncertainty indicator.
@@ II-C Temporal Pooling Layers
Temporal pooling is a critical component in speaker embedding extraction, as it aggregates frame-level features into utterance-level representations [2, 57]. Early approaches rely on statistical pooling, which summarizes the entire frame sequence using global statistics. Temporal Average Pooling (TAP) computes the mean of frame-level features, while Temporal Statistics Pooling (TSTP) and Temporal Standard Deviation Pooling (TSDP) further incorporate variance information to enhance robustness [8, 9, 10].
A key limitation of statistical pooling is the assumption that all frames contribute equally to the utterance-level representation. To address this issue, attention-based pooling methods assign different weights to frames according to their speaker-discriminative importance. Attentive Statistics Pooling (ASP) computes weighted mean and standard deviation vectors, and subsequent studies show that multi-head attention further improves pooling performance [11, 12, 13, 14, 15]. ECAPA-TDNN extends this idea by introducing channel-wise attention, enabling the model to emphasize speaker characteristics across both temporal and channel dimensions [16].
Instead of explicitly learning attention weights, the xi-vector model [24] introduces uncertainty-aware pooling, where frame-level uncertainty estimates are used to modulate each frame’s contribution to the utterance-level representation. As adopted in this work, this formulation provides a principled and interpretable mechanism for down-weighting unreliable frames while preserving speaker-discriminative information.
@@ III-A Uncertainty-aware Speaker Model: xi-vector
$\boldsymbol{\phi}^{\text{s}}=\left(\frac{\boldsymbol{\phi}-\boldsymbol{\mu}_{\text{bn}}}{\sqrt{\boldsymbol{\sigma}_{\text{bn}}+\epsilon\mathbf{I}}}\otimes\boldsymbol{\gamma}_{\text{bn}}+\boldsymbol{\beta}_{\text{bn}}\right)\mathbf{A}_{\text{fc}}^{\top}+\mathbf{b}_{\text{fc}},$ (5)
@@
$\mathbf{\Sigma}^{\text{s}}=\mathbf{A}_{\text{fc}}\frac{\mathbf{L}^{-1}\otimes\boldsymbol{\gamma}_{\text{bn}}^{2}}{\boldsymbol{\sigma}_{\text{bn}}+\epsilon\mathbf{I}}\mathbf{A}_{\text{fc}}^{\top}.$ (6)
@@ III-B Softmax in Speaker Classification
As shown in Fig. 1, the projection layer maps the utterance-level representation $\boldsymbol{\phi}^{\text{s}}$ into a $C$-dimensional vector, where $C$ is the number of speakers in the training set. This vector is then optimized using a classification objective, with the softmax function playing a central role.
@@ III-B2 Additive Angular Margin Softmax
The conventional Softmax loss does not explicitly enforce intra-class compactness or inter-class separability of the learned embeddings. To address this limitation, Additive Angular Margin Softmax (AAM-Softmax) [43] enhances discriminative power by introducing an angular margin in the cosine similarity space.
Specifically, AAM-Softmax removes the bias term and applies $\ell_{2}$ normalization to both the class weights and speaker embeddings, such that the classification decision depends solely on their cosine similarity. An additive angular margin $m$ is then imposed on the target class, yielding the following loss:
$\mathcal{L}_{\text{AAM}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{s\cos(\theta_{y_{i}}+m)}}{e^{s\cos(\theta_{y_{i}}+m)}+\sum_{j\neq y_{i}}e^{s\cos\theta_{j}}},$ (8)
where $\theta_{j}$ denotes the angle between the normalized embedding and the $j$-th class weight, $s$ is a scaling factor, and $m$ controls the angular margin.
@@ IV-A Stochastic Variance Loss
Our previous work xi-vector [24] relies solely on a classification loss to implicitly supervise frame-level uncertainty, without any form of explicit supervision. Such implicit uncertainty learning often results in suboptimal or unstable estimation. To overcome this limitation, we introduce an explicit supervision mechanism for utterance-level uncertainty modeling, termed Stochastic Variance Loss (SVL). Our initial work [36] firstly proposed SVL in the context of speaker recognition and demonstrated its effectiveness. In this paper, we further generalize this idea by exploring other alternative formulations to realize the core concept of SVL.
The proposed loss $\mathcal{L}_{\text{SVL}}$ aims to establish a pseudo ground truth for utterance-level uncertainty estimation. Specifically, we treat the distance between the centroid of embeddings from the same speaker and the current utterance’s embedding $\boldsymbol{\phi}^{\text{s}}$ as a pseudo label.
Suppose we use $\boldsymbol{\phi}^{\text{c}}_{y_{i}}$ to represent the centroid of speaker $y_{i}$ of utterance $i$:
$\boldsymbol{\phi}^{\text{c}}_{y_{i}}=\frac{\sum_{i=1}^{I}\boldsymbol{\phi}^{\text{s}}_{i}}{I},$ (9)
where all $I$ utterances belong to the same speaker $y_{i}$. The loss $\mathcal{L}_{\text{SVL}}$ is given by:
$\mathcal{L}_{\text{SVL}}=\frac{1}{B}\sum_{b=1}^{B}\|\alpha\mathbf{(\Sigma}^{\text{s}}_{b})^{\frac{1}{2}}-|\boldsymbol{\phi}^{\text{s}}_{b}-\boldsymbol{\phi}^{\text{c}}_{y_{b}}|\|^{2},$ (10)
where $B$ denotes the number of batch size in training. The matrices of current batch $\mathbf{\Sigma}^{\text{s}}_{b}$ and $\boldsymbol{\phi}^{\text{c}}_{y_{b}}$ denote the covariance matrix for utterance $b$ and the centroid of speaker $y_{b}$, respectively. The symbol $\|\cdot\|$ denotes the Euclidean norm. In addition, $\alpha$ is a learnable parameter initialized to $1$, referred to as the uncertainty scaling factor. It aimes to address potential scale mismatch between predicted uncertainty and pseudo ground truth. Since only the embedding distance of a single utterance-centroid pair is used to approximate the uncertainty (i.e., variance), we refer to this approach as stochastic variance learning.
@@
  • Pre: The centroid is computed using embeddings $\boldsymbol{\phi}^{\text{s}}$ extracted from a pre-trained model with the same architecture, but was trained solely with the cross-entropy loss $\mathcal{L}_{\text{AAM}}$. As a result, the speaker centroids remain fixed throughout training and are subsequently used to provide supervision, as defined in  (10).
  • Epoch: Instead of relying on a pre-trained model, the speaker embedding $\boldsymbol{\phi}^{\text{s}}$ is extracted using the checkpoint from the last epoch. Then centroids are computed after each training epoch. This allows the centroids to evolve along with model optimization, leading to progressively more accurate estimate.
@@
where $\kappa$ is used to control loss weight and is dynamically adjusted based on the current epoch $\psi$ to ensure that $\mathcal{L}_{\text{SVL}}$ gradually contributes as training progresses:
$\kappa=\begin{cases}0,&\text{if }\text{epoch}\leq\psi_{\text{SVL}},\\ \lambda\frac{\psi-\psi_{\text{SVL}}}{\psi_{\text{Max}}-\psi_{\text{SVL}}},&\text{otherwise},\end{cases}$ (12)
where $\lambda$ is a pre-defined fixed constant, $\psi_{\text{SVL}}$ indicates epoch when $\mathcal{L}_{\text{SVL}}$ starts to be applied. $\psi_{\text{Max}}$ indicates the maximum number of training epochs. In this paper, we set $\psi_{\text{SVL}}=40$ and $\psi_{\text{Max}}=150$. Cause we use WeSpeaker [58, 59] as training tool, its margin changes before $40$ epoch, which causes unstable training and produces NaN if we incorporate $\mathcal{L}_{\text{SVL}}$ at this stage.
@@ IV-B Uncertainty-aware Scoring
$s_{u}=\frac{<\boldsymbol{\phi}^{\text{s}}_{2},\boldsymbol{\phi}^{\text{s}}_{1}>}{\sqrt{\boldsymbol{\phi}{{}^{s}_{2}}^{\top}(\mathbf{I}+\rho\boldsymbol{\Sigma}^{\text{s}}_{2})^{-1}\boldsymbol{\phi}^{\text{s}}_{2}}\sqrt{\boldsymbol{{\phi}}{{}^{s}_{1}}^{\top}(\mathbf{I}+\rho\boldsymbol{\Sigma}^{\text{s}}_{1})^{-1}\boldsymbol{\phi}^{\text{s}}_{1}}},$ (13)
where $\rho$ denotes a scaling factor in cosine scoring, and $\mathbf{I}$ is the identity matrix. When the covariance matrix $\boldsymbol{\Sigma}^{s}$ or $\rho$ is zero, the uncertainty-aware cosine scoring $s_{u}$ reduces to the normal cosine similarity. Wang et al. [38] have proposed to use the fixed constant $1/d$ as $\rho$, where $d$ denotes the dimension of speaker embedding. However, a fixed value may not be optimal, as demonstrated in our previous work, xi+ [36]. To address this limitation, we introduce a learnable uncertainty scaling factor $\alpha$, which replaces $\rho$ in uncertainty-aware scoring. This parameter can be adapted to different training datasets, allowing the model to optimally calibrate the influence of uncertainty in the scoring process.